If angular displacement of a wheel moving on a curved path is Θ = 1.5t + 2t2 , where t is in seconds. Radius of wheel is 60 cm. Angular velocity and linear velocity at t = 2sec. will be
 
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
angular velocity = dΘ/dt = 1.5 + 4t
at t=2 sec
angular velocity = 1.5 + 4*2 = 9.5 rad/sec
radius = 0.6 m
linear velocity = (angular velocity)*radius = 9.5*0.6 = 5.7 m/s
Hope it clears.............................